### v1

# Learning with less: label-efficient land cover classification at very high spatial resolution using self-supervised deep learning

Abstract
Deep learning semantic segmentation methods have shown promising performance for very high resolution land cover classification, but the challenge of collecting large volumes of representative training data creates a significant barrier to widespread adoption of such models for meter-scale land cover mapping over large areas. In this study, we present a novel label-efficient approach for statewide land cover classification using only 1,000 annotated reference image patches with self-supervised deep learning. We use the “Bootstrap Your Own Latent” pre-training strategy with a large amount of unlabeled color-infrared aerial images (377,921 256256 pixel patches) to pre-train a ResNet-101 convolutional encoder. The learned encoder weights were subsequently transferred into multiple deep semantic segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN), which were then fine-tuned using very small training dataset sizes with cross-validation (250, 500, 750 patches). Among the fine-tuned models, we obtained the 87.14% overall accuracy and 75.58% macro F1 score using an ensemble of the best performing U-Net models for comprehensive , 8-class land cover mapping, covering more than 123 billion pixels over the state of Mississippi, USA. Detailed qualitative and quantitative analysis revealed accurate mapping of open water and forested areas, while highlighting challenges in accurate delineation between cropland, herbaceous, and barren land cover types. These results show that self-supervised learning is an effective strategy for reducing the need for large volumes of manually annotated data, directly addressing a major limitation to high spatial resolution land cover mapping at scale.
keywords:
[…]
1 Introduction
Land cover data products provide valuable information about the characteristics of the Earth’s surface for a wide variety of applications, including agriculture, urban planning, and environmental monitoring [1, 2, 3, 4, 5]. Research into automated methods for land cover classification from remotely sensed data has been ongoing for several decades, with a majority of methods relying on supervised learning techniques to classify land cover using multispectral aerial or satellite imagery [6, 7, 8, 9, 10]. Historically, pixel-wise and object-based image analysis (OBIA) classification methods have been the predominant approach for land cover classification, where spatial units (i.e., pixels or objects) in the image are classified based on their spectral reflectance, hand-crafted features, and ancillary data from other geospatial datasets [11, 12]. Very high spatial resolution (VHSR, ) land cover products offer detailed information about the Earth’s surface with high spatial accuracy, which is valuable for applications that require fine-grained information about land cover types and their spatial distribution [13, 14]. However, the creation of such VHSR data products is inherently challenging due to the complexity of the land cover patterns at these scales and the increased spectral variability within classes [15, 16, 17]. In this context, the application of deep convolutional neural networks (CNNs) to land cover classification tasks has gained significant attention in recent years, as these models excel at extracting spatial and spectral features from imagery in a data-driven manner, making them well-suited for VHSR land cover classification tasks [18, 19, 20]. Nevertheless, these models require far more labeled training data compared to traditional machine learning methods, serving as a significant barrier to their widespread adoption for operational land cover mapping at VHSR [21, 22].
This need for large amounts of ground-truth data for accurate training of deep learning models has led to extensive research into methods for reducing the amount of data necessary to train these models [23, 24, 25]. Most of these methods leverage transfer learning procedures, where the parameters of a model pre-trained using a large dataset are used to initialize a model that is subsequently fine-tuned on a smaller dataset that is more relevant to the target task [26, 27]. For example, a common approach to transfer learning for land cover classification tasks is to pre-train a model on a large general-purpose image classification dataset (e.g., ImageNet [28]) and then fine-tune the model on a smaller dataset of interest [29, 30, 31, 32]. Despite being effective, the domain shift between the source and target datasets can hinder the model’s performance, and the extent of this domain shift between general-purpose image classification datasets and VHSR remote sensing datasets is not well understood [33]. As fully-supervised pre-training requires large-scale labeled datasets, which are relatively scarce in the remote sensing domain for VHSR imagery, there is a need for alternative pre-training strategies that reduce the reliance on labeled data [34, 35].
More recently, self-supervised pre-training methods have been proposed as an alternative to traditional fully-supervised pre-training methods, enabling models to learn useful spatial and spectral information from input imagery through the use of pretext tasks that do not require labeled data [36, 37, 38]. These techniques are particularly attractive for remote sensing applications, as large amounts of labeled imagery datasets are often unavailable, but source imagery is widely available due to the growing number of Earth observation satellites and aerial imagery platforms [39]. However, few studies have rigorously investigated the applicability of these methods to operational land cover mapping at VHSR, especially in scenarios where ground-truth data are scarce. Most research evaluating self-supervised models pre-trained with remote sensing data uses benchmark VHSR land cover datasets as a proxy to evaluate the quality of pre-trained models, with the ISPRS Potsdam [40] and Vaihingen [41] datasets seeing wide use for this purpose [42, 43, 44, 45]. However, these datasets contain relatively large amounts of labeled training data, and thus may not necessarily be representative of real-world scenarios where no available VHSR land cover data exists and costly manual annotation is required to create such data (e.g., by experience, a single 256256 image patch can take upwards of 30 minutes to label well). Furthermore, while there is growing momentum in the development of large remote sensing foundation models using self-supervised techniques [46, 39], there is still a critical need to systematically assess how these approaches can enhance VHSR land cover mapping performance in practical contexts, a gap this study directly addresses.
Self-supervised approaches inspired by masked language modeling in natural language processing have been shown to be strong approaches for learning useful feature representations from images without the need for labeled data [47, 48]. Masked image modeling (MIM) methods take advantage of the Vision Transformer (ViT) patching mechanism to randomly mask out a portion of the input image and then train the model to reconstruct the missing patches from the visible patches [49, 50]. MIM techniques have become the predominant self-supervised pre-training paradigm for creating remote sensing foundation models, which are large pre-trained models that can be adapted for a wide variety of tasks and domains that ideally outperform models trained from scratch or use more traditional transfer learning approaches [51, 45]. However, while powerful, ViTs lack the inductive biases toward imagery that make CNNs effective at extracting spatial features from images [52]. This is problematic, as ViTs typically only begin to outperform CNNs when pre-trained on very large datasets [53], a practical consideration for researchers and practitioners who must balance both computational resources and data availability when developing models. Self-distillation techniques, such as Bootstrap Your Own Latent (BYOL) [54] and DINO [55], are alternatives to MIM pre-training that can be used to learn robust and transferable feature representations without the need for negative samples, large batch sizes, or ViT backbones, which are hard requirements for many other self-supervised learning methods [38, 56]. Such approaches largely revolve around maximizing the similarity between feature representations of different augmented views of the same input using a teacher-student framework [57]. While the student network is trained directly using standard backpropagation and gradient descent algorithms, the teacher’s parameters are updated using a momentum-based rule derived from the student’s parameters. Self-distillation methods are attractive for remote sensing applications due to their inherent flexibility and ability to learn strong feature representations using a variety of backbone architectures, including CNNs [58]. Given the architectural constraints of MIM and the computational demands of contrastive methods, self-distillation techniques offer a compelling, architecture-agnostic solution, suitable for pre-training CNNs on domain-specific, moderate-sized datasets [59].
In this study, we propose an innovative label-efficient framework that addresses the challenge of operational land cover mapping at resolution under label-scarce scenarios through self-supervised learning. Specifically, our framework used BYOL as a pretext task during pre-training, enabling a ResNet-101 convolutional encoder network to learn to extract deep features from unlabeled 377,921 256256 patches of color-infrared aerial imagery without the need for full supervision using labeled data. We then transferred the learned encoder weights into multiple semantic segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN) that are subsequently fine-tuned using very small training dataset sizes (250, 500, and 750 256256 patches) using cross-validation. We classified 8 classes of land cover (open water, impervious structures, impervious surfaces, barren land, forest/woody vegetation, herbaceous/low vegetation, cultivated crops, and unclassified (i.e., shadow)) over the entire state of Mississippi, USA using an ensemble of the best-performing models, totaling over 123 billion pixels. Validation and accuracy assessment were performed using a spatially independent test set of 25,000 randomly sampled points with manually verified labels to ensure robust and unbiased evaluation of our approach and the resulting land cover map. We demonstrated that these approaches can be used to train deep semantic segmentation models that are accurate, generalizable, and capable of yielding reasonable performance even under extreme low-data scenarios, even when only 3 bands (red, green, and near-infrared) are available in the input imagery. Our ultimate aim is to reduce practical barriers to producing high spatial resolution land cover maps while contributing to ongoing efforts to evaluate and contextualize the implications of self-supervised learning techniques for remote sensing purposes.
2 Data sources
2.1 NAIP aerial imagery
The USDA’s National Agricultural Imagery Program (NAIP) is a U.S. federal project that provides a publicly available archive of high-resolution multispectral imagery over the conterminous United States, with a spatial resolution of , a four-band spectral resolution (near-infrared, red, green, and blue), delivered as 8-bit orthorectified tiles [60]. NAIP imagery is collected on a state-by-state basis during the agricultural growing season, and each state is imaged once every 2–3 years. The lack of use restrictions on NAIP imagery has made it a popular choice for high spatial resolution remote sensing applications in the United States, including building footprint extraction [61], forest measurements [62, 63], and land cover classification [64, 65, 66, 67]. Color-infrared (CIR) composites are generated using the near-infrared, red, and green bands to emphasize vegetation features in the imagery. All 2023 CIR composites were obtained as county-level composites from the USDA NRCS Geospatial Data Gateway [68] in MrSID format in the appropriate UTM zone for each county (UTM Zone 15N for Western Mississippi, UTM Zone 16N for Eastern Mississippi) at resolution. Each county-level composite was converted to GeoTiff format, reprojected to the Mississippi Transverse Mercator coordinate reference system (EPSG:3813), and resampled to resolution using bilinear interpolation. The total size of the NAIP imagery used in this work was , comprising 82 county-level GeoTiff files, an average of per county, with individual county rasters ranging in size from to . A visualization of the NAIP imagery used over the study area is provided in Figure 1. We selected three-band images for compatibility with BYOL, which is designed for three-channel inputs, and for easy comparison with models using ImageNet pre-trained weights, which also expect three-channel inputs.
2.2 NLCD land cover
The stratification approach proposed in this work (Section 3.1) relies on an existing land cover product to select strata from which training samples are drawn. For this purpose, we used the USGS’s National Land Cover Database (NLCD) to provide the necessary land cover information for stratification. The NLCD has historically been among the most widely used land cover datasets in the United States, supporting numerous applications in human and ecosystem health studies, urban planning, natural hazard assessment, hydrology, and land management [69]. The latest version of the NLCD (Annual NLCD) provides a yearly 20-class land cover product primarily derived from Landsat imagery at resolution from 1985 to 2024 [70]. While the spatial resolution of the NLCD land cover product is substantially coarser than the land cover product we aim to produce, the NLCD provides a comprehensive land cover product at a national scale and has a history of consistent use in the literature [71, 72, 73], making it a suitable choice to provide land cover information to guide the stratification process. To align with the collection year of the NAIP imagery used in this work, we use 2023 data over the state of Mississippi.
3 Methodology
Our proposed methodology follows a multi-stage approach to model development that investigated the implications of self-supervised pre-training with the goal of producing an accurate VHSR land cover product over the state of Mississippi using a limited amount of annotated training data (Figure 2). In Section 3.1, we describe a sampling strategy designed to capture the diversity of land cover across Mississippi. Candidate patches were selected and grouped based on their land cover composition, and a subset was sampled for dense visual annotation. A point-based assessment dataset of 25,000 samples was also created for model evaluation and final product accuracy assessment. Section 3.2 describes how BYOL was used to pre-train a ResNet-101 CNN using 377,921 patches of unannotated pre-training data from the area of interest (excluding the already sampled regions). Section 3.3 describes the fine-tuning and evaluation process, where the pre-trained ResNet-101 backbone was integrated into various semantic segmentation architectures and fine-tuned on limited annotated training data using stratified k-fold cross-validation. We evaluated the performance of multiple segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN) using varying amounts of training data (250, 500, 750 samples) to understand the implications of model architecture and training dataset size on final model performance. Finally, several fine-tuned models were compiled into a model ensemble to produce a land cover map at resolution over the state of Mississippi for the year 2023, with the final data product undergoing a thematic accuracy assessment using the point-based assessment dataset, as described in Section 3.4.
3.1 Stratified sampling and high-quality land cover training dataset generation
Stratifying the sampling of training data is a common practice in remote sensing to ensure that the training dataset is representative of the land cover distribution of the area of interest, particularly when the area contains rare or underrepresented classes [74, 75, 76]. To ensure representative sampling for k-fold cross-validation, we employed a stratified sampling strategy to select 1,000 candidate training samples for annotation. We overlaid a grid of 256256 pixels (i.e., 256 patches), forming a population of 1,889,606 candidate patches used to support both the unlabeled dataset for pre-training and the labeled dataset for fine-tuning. To curate the fine-tuning dataset, zonal histograms of land cover class distribution were extracted from the 2023 NLCD dataset for each candidate patch to characterize its land cover composition as a type of feature vector. We then applied PCA to reduce the dimensionality of the feature vectors and K-means clustering to form 250 strata based on land cover composition. Finally, we randomly sampled 4 candidate training samples from each stratum, placing each into a separate fold of the k-fold training dataset, resulting in 4 folds totaling 1,000 samples. Figure 3 shows the NLCD land cover across Mississippi, along with the spatial distribution of the selected samples and the class composition within each fold.
A team of trained annotators visually annotated the selected 1,000 samples according to the land cover classification scheme in Table 1. Samples were randomly assigned to each annotator to ensure that each fold contained a mix of contributions from different individuals. After the annotators completed their work, the annotations were reviewed and corrected by another annotator whose task was to ensure that the annotations were consistent and accurate across all samples. Annotators employed the Segment Anything Model 2 (SAM 2) hosted on the Computer Vision Annotation Tool (CVAT) annotation platform to assist with the visual annotation process [77, 78] by proposing initial object masks that annotators could manually refine as needed. Examples of annotated samples are shown in Figure 4.
| Index | Class | |
| 1 | Open Water | |
| 2 | Impervious Structures | |
| 3 | Impervious Surfaces | |
| 4 | Barren Land | |
| 5 | Forest/Woody Vegetation | |
| 6 | Herbaceous/Low Vegetation | |
| 7 | Cultivated Crops | |
| 8 | Unclassified |
In addition to the 1,000 samples selected for annotation, a large dataset of unannotated samples was also collected from the state of Mississippi using simple random sampling from the remaining candidate patches to form a large dataset of unannotated samples for pre-training the ResNet-101 encoder. 20% (377,921) of the total number of candidate samples were selected for pre-training, while 5% (94,480) were selected to serve as a pre-training validation dataset to monitor for overfitting during the pre-training process. Once all sample patches were selected, an additional point-based assessment dataset of 25,000 samples was created by randomly sampling points from the remaining candidate samples to serve as a testing dataset for the purposes of model evaluation and final accuracy assessment of the resulting land cover product. No points within of a sampled patch were selected to prevent data leakage caused by spatial autocorrelation, and no points within of a previously sampled point were selected to ensure that the assessment points were spatially independent of each other. These points were visually annotated by a single annotator using 2023 NAIP imagery. For model evaluation purposes, 256 patches were extracted around each point such that the point was at the center of the patch, and a model’s prediction for the central pixel corresponding to the point was compared to the point label to compute accuracy metrics.
3.2 Self-supervised pre-training using BYOL
The pre-training phase of the methodology leverages the large dataset (377,921) of unannotated samples such that a ResNet-101 encoder can be trained to extract relevant features from the imagery that are well-suited for downstream land cover classification tasks. BYOL uses a “self-distillation” approach to learn representations from images using two separate but interdependent networks: an online network containing the image encoder, a projection head, and a prediction head, and a target network that is structurally similar to the online network but lacks the prediction head, initialized using a copy of the online network’s parameters at the start of training [54]. Given an input image, two separate augmented views of the image are created using a set of transformations. As illustrated in Figure 5, each view is then passed through the online and target networks to produce feature vectors. The online network, which processes view , is then trained to predict the feature vector produced by the target network from view . However, the target network is not trained directly; instead, it is updated using an EMA of the online network weights, giving the online network a stable yet constantly updated target to learn from. Ultimately, the image encoder from the online network is retained after training to be transferred to downstream tasks.
The online encoder’s parameters are updated with respect to the loss function in Equation 1, where is the output of the online network and is the output of the target network.
The target network’s parameters are updated according to the EMA update rule shown in Equation 2, where is the current set of parameters for the online network and is the decay factor for the EMA operation, and is increased from 0.996 to 1 over the course of pre-training according to a cosine annealing schedule.
We applied BYOL to a ResNet-101 encoder network using the dataset of 377,921 unannotated samples from the state of Mississippi. We tested two configurations of the BYOL pre-training process: one starting with ImageNet pre-trained weights at the beginning of the pre-training process and one starting with randomly initialized weights. The pre-training process was run for 300 epochs, with the learning rate reduced from to 0 over the course of training using a cosine annealing schedule, with a warmup period of 10 epochs to linearly increase the learning rate from 0 to . The AdamW optimizer was used with a batch size of 4096: to achieve this large batch size while working within the memory constraints of a single Nvidia A100 GPU with 80 GB of VRAM, we employed gradient accumulation using a microbatch size of 256. After each epoch, the model was evaluated on the pre-training validation dataset using the BYOL loss function (as during training) to monitor for overfitting.
To evaluate the quality of the learned representations, we used a linear probing approach to determine how well the deep features extracted by the pre-trained networks could be used to classify land cover classes using a simple linear classifier. To accomplish this, the feature maps produced by the ResNet-101 encoder were upsampled to the original input resolution using bilinear interpolation, and a linear classifier was trained on top of these feature maps using the annotated training dataset from Section 3.1. The linear classifier consisted of a single 11 convolutional layer with a softmax activation function to produce class probabilities for each pixel. K-fold cross-validation was employed to minimize the influence of variation in the training dataset while evaluating the performance of this method across different training dataset sizes (); the splitting strategy for each is shown in Figure 6.
A separate model was trained for each run of the cross-validation framework, with the model being trained for a maximum of 1,000 epochs using the AdamW optimizer with a learning rate of and a batch size of 32 [79]. Focal loss was used as the loss function to address the class imbalance in the land cover classes as shown in Equation 3, where is the model’s estimated probability for the true class and [80].
A warmup period of 10 epochs was used to linearly increase the learning rate from to at epoch 10, after which the learning rate was decayed by a factor of 0.1 when the validation loss failed to decrease for 10 epochs. After 50 epochs of no improvement in the validation loss, training was halted and the model with the lowest validation loss was selected as the final model for that run. Automatic mixed precision (AMP) training was used to speed up the training stage and reduce memory usage. Due to the limited size of the training datasets, we employed a strong set of image augmentations on-the-fly to ensure that the models did not overfit to the training data, including random horizontal and vertical flips (50% probability each), random 90∘ rotations (0∘, 90∘, 180∘, 270∘; 100%), color jitter (brightness = 0.4, contrast = 0.4, saturation = 0.2, hue = 0.2; 80%), random grayscale conversion (20%), and Gaussian blur (kernel size = 25, ; 100%). Since each cross-validation run used its validation split to tune the optimizer and trigger early stopping, we did not use validation-split metrics for final performance reporting. Rather, the performance of the linear probing models was evaluated using the point-based assessment dataset (25,000 samples), with the results being compared between the two BYOL pre-training configurations (i.e., different weight initialization strategies) as well as a baseline model that used the ResNet-101 encoder with ImageNet pre-trained weights without any BYOL pre-training. Considering the natural imbalance of the land cover classes, we chose the macro F1 score as the primary evaluation metric, as shown in Equation 4, where is the number of classes, is the number of true positives for class , is the number of false positives for class , and is the number of false negatives for class .
3.3 Implementation of multiple semantic segmentation architectures with fine-tuning under limited labeled dataset
Once the ResNet-101 encoders were pre-trained using BYOL, they were used as a backbone for a semantic segmentation architecture that was fine-tuned using the annotated training dataset from Section 3.1. We evaluated 6 different semantic segmentation architectures to determine which framework was best suited for the land cover classification task at hand. We excluded any architectures that rely on transformer-based encoders (e.g., SegFormer [81], SETR [82]) or specialized convolutional backbones (e.g., HRNet [83], BiSeNetV2 [84]) to maintain a controlled comparison between architectures that all use the same ResNet-101 encoder network. Table 2 shows the semantic segmentation architectures that were evaluated in this study, and a brief description of each architecture is provided below.
| Model | Number of Parameters | |
| FCN | 42,528,856 | [85] |
| U-Net | 92,302,024 | [86] |
| Attention U-Net | 83,610,272 | [87] |
| DeepLabV3+ | 58,199,128 | [88] |
| UPerNet | 73,251,400 | [89] |
| PAN | 140,688,970 | [90] |
FCN
The Fully Convolutional Network (FCN) was one of the earliest deep learning architectures designed specifically for semantic segmentation tasks [85]. As its name suggests, FCN is composed entirely of convolutional layers, eschewing the use of fully connected layers that are common in image classification architectures. Under this design, instead of outputting a single class label for an entire image, the model outputs a spatial map of class probabilities that correspond to the input image’s spatial dimensions. FCN uses a standard convolutional encoder to extract hierarchical features from the input image, followed by a series of upsampling layers to produce the final segmentation map. Though simple, FCN was a groundbreaking architecture that demonstrated the potential of deep learning for semantic segmentation tasks, and its encoder-decoder framework has been widely adopted in subsequent architectures. FCN saw early success in remote sensing applications, particularly for land cover classification tasks at high spatial resolutions [91, 92, 93]. We implemented the FCN-8s variant of FCN, which uses skip connections to combine feature maps from the encoder at multiple stages to produce more accurate segmentation maps.
U-Net
U-Net is a popular semantic segmentation architecture that was among the early adopters of the encoder-decoder framework for semantic segmentation tasks. Under such a framework, the encoder network extracts hierarchical features from the input image, while the decoder network learns a mapping from these low-dimensional features to an output segmentation map that matches the original input resolution [86]. Notably, U-Net’s decoder network is approximately symmetric to the encoder network, meaning that the feature maps are upsampled in stages that mirror the downsampling stages of the encoder. As such, skip connections can be used to pass feature maps from the encoder to the decoder at corresponding stages: the encoder’s feature maps are concatenated with the decoder’s feature maps at each stage to provide the decoder with relevant, resolution-specific information that otherwise would be lost during the downsampling process. U-Net’s simple yet powerful architecture has made it a popular choice for a wide variety of remote sensing applications [94, 95, 96, 97] Our U-Net implementation differed from the original U-Net architecture in two key ways: first, we opted to use simple bilinear interpolation for upsampling in the decoder rather than transposed convolutions to reduce the presence of checkerboard artifacts in the output segmentation maps (see [98] for more details); second, as we used a ResNet-101 encoder that downsamples the input image by a factor of 4 in the first stage (as opposed to the factor of 2 downsampling in the original U-Net), we modified the decoder to account for this difference by adding an additional upsampling stage to ensure that the output segmentation map matches the input resolution.
Attention U-Net
As the name suggests, Attention U-Net is a variant of the U-Net architecture that incorporates attention mechanisms that emphasize spatial regions of interest in the feature maps. This is accomplished through the use of attention gates along the skip connections between the encoder and decoder, which learn to weight the features passed from the encoder to the decoder based on their relevance to the current task. By focusing on the most relevant spatial regions, Oktay et al. [87] demonstrated that Attention U-Net achieved improved performance on medical image segmentation tasks compared to the original U-Net architecture, particularly with regard to the localization of object boundaries. Like U-Net, our Attention U-Net implementation used bilinear interpolation for upsampling in the decoder rather than transposed convolutions, and the decoder included an additional upsampling stage to account for the ResNet-101 encoder’s downsampling factor of 4 in the first stage.
DeepLabV3+
DeepLabV3+ is an evolution of the DeepLab family of semantic segmentation architectures that incorporates atrous (dilated) convolutions and spatial pyramid pooling to capture multi-scale context in the feature maps [88]. At the core of DeepLabV3+ is an Atrous Spatial Pyramid Pooling (ASPP) module that aggregates multi-scale context from the final feature maps produced by the encoder. This ASPP module uses parallel atrous convolutions to encode feature maps at multiple scales as opposed to using pooling operations: the use of atrous (i.e., dilated) convolutions allows the network to capture features at multiple scales in a learned manner. A comparatively simple decoder network is utilized that upsamples the output of the ASPP module using bilinear interpolation, concatenates it with low-level feature maps from the encoder, and upsamples the result to produce the final segmentation map. Like U-Net, DeepLabV3+ has seen wide adoption for remote sensing applications [99, 100, 101] DeepLabV3+ was originally designed to use a modified ResNet-101 encoder, and as such minimal effort was necessary to replace the original encoder with our pre-trained ResNet-101 encoder.
UPerNet
Unlike most semantic segmentation architectures that only attempt to classify objects at a single scale, UPerNet (Unified Perceptual Parsing Network) is designed to classify objects as part of a hierarchical structure that operates at multiple levels of granularity. UPerNet classifies an image at one global level (i.e., scene classification) and at four local levels (object, part, material, and texture classification maps) [89]. The core component of UPerNet is Feature Pyramid Network (FPN), a U-Net-like encoder-decoder architecture that serves as the primary feature extractor for UPerNet. The encoder network’s downsampling stages are similar to those of ResNet-101, making substituting the original FPN encoder with a ResNet-101 encoder straightforward. A pyramid pooling module is placed after the encoder network in the FPN to aggregate global context from the feature maps prior to passing them to the decoder. Our implementation disregarded the multi-level classification aspect of UPerNet, and only used the object-level classification head to produce a single land cover classification map. This object-level classification head receives feature maps from all levels of the FPN decoder, classifies each pixel, and upsamples the output to match the input resolution. Classifying land cover using feature maps from all levels of the decoder allows the classification head to leverage features at multiple scales to inform its predictions, unlike more traditional encoder-decoder architectures that only consider the output from the final, highest-resolution feature map of the decoder.
PAN
Incorporating multi-scale context into semantic segmentation models is a common theme among many modern architectures, but many of these frameworks utilize sub-networks or modules that are limited in their ability to capture highly complex relationships in the feature maps and re-weight them accordingly to emphasize the most relevant features. PAN (Pyramid Attention Network) attempts to address this limitation through the use of a Feature Pyramid Attention (FPA) network module that utilizes convolutions with multiple kernel sizes to perform a spatial attention operation on the feature maps produced by the encoder [90]. Like the ASPP module in DeepLabV3+, PAN uses this FPA module with the goal of refining large-scale context in the final feature maps produced by the encoder. However, PAN’s FPA module aims to retain localization information in the feature maps through the use of convolutional attention rather than pooling or atrous convolutions. Similar to Attention U-Net, Global Attention Upsample (GAU) modules are used in the decoder to serve as a form of gating mechanism that attends to the most relevant channels in the feature maps before adding them to the upsampled feature maps from the previous stage of the decoder (as opposed to attending spatially and concatenating the feature maps in Attention U-Net). ResNet-101 is used as the encoder in the original PAN architecture, and as such no modifications were necessary to incorporate our pre-trained ResNet-101 encoder.
Each model was fine-tuned using the same configuration as the models in Section 3.2 (i.e., AdamW optimizer, learning rate of , batch size of 32, focal loss function, AMP, etc.). During fine-tuning, the weights of the ResNet-101 encoder were unfrozen and updated along with the weights of the semantic segmentation architecture. Evaluation was again performed using cross-validation across 250, 500, and 750 training samples (Figure 6), and the performance of each model was assessed using each model’s macro F1 score on the point-based assessment dataset (25,000 samples) to select the best-performing combination of pre-training procedure, semantic segmentation architecture, and training dataset size using the mean macro F1 score across the k-folds.
3.4 Model ensembling, large-scale land cover mapping, and validation
Once the best-performing combination of pre-training procedure, semantic segmentation architecture, and training dataset size was determined, we used the models trained using this configuration across all folds to produce a model ensemble for large-scale land cover mapping over the state of Mississippi. All models in the ensemble shared the same architecture, pre-training procedure, and training dataset size, differing only in the specific samples used for training due to the k-fold cross-validation procedure. For example, if the best-performing configuration is found to be a U-Net architecture with BYOL pre-training using 750 training samples, then the final ensemble would consist of 4 U-Net models, each pre-trained using BYOL and fine-tuned using 750 training samples from each fold of the cross-validation procedure, with each model having seen a different subset of the training data. Model ensembling was chosen as it is a popular approach to improving the robustness and accuracy of machine learning models by combining the predictions of multiple distinct models to produce a final prediction. Each model in the ensemble produced a land cover probability vector for each pixel in the input image, and these class probability vectors were averaged across all models in the ensemble to produce a final class probability vector for each pixel, with the final predicted class being the class with the highest probability from the averaged class probabilities [102]. A “sliding window” approach was used to process the imagery in patches with a stride of to ensure that each pixel would be seen multiple times to further maximize the effects of the model ensemble. Additionally, test-time augmentations were employed to further improve the robustness of the predictions: at any given location, each model in the ensemble received an input that was randomly augmented using random horizontal and vertical flips along with random 90∘ rotation [103]. For example, if there are 4 models in the ensemble, each model would receive a different random augmentation of the input patch, resulting in 4 different predictions for that patch. The final output patches of the ensemble were weighted using element-wise multiplication with a 256256 Gaussian kernel to de-emphasize the probabilities of pixels near the edge of the classification maps in order to mitigate presence of edge artifacts in the final merged output. These inference steps were used to classify the entire state of Mississippi using 2023 NAIP imagery at resolution, amounting to over 123 billion pixels. A thematic accuracy assessment of the final land cover product was performed using the point-based assessment dataset (see Section 3.1). Overall accuracy, per-class user accuracy, and per-class producer accuracy metrics were computed to assess the quality of the final product, as shown in Equations 5, 6, and 7, where TP is the number of true positives, FP is the number of false positives, and FN is the number of false negatives.
4 Results
The evaluation of the self-supervised learning approach is divided into two parts: linear probing performance and fine-tuning results. The linear probing experiments are intended to provide a measure of the quality of the learned image representations produced by the pre-trained encoder in isolation (i.e., without any deep segmentation heads or fine-tuning). As shown in Figure 7, using the embeddings from a ResNet-101 model pre-trained using BYOL as input to a linear classifier yielded substantial improvements in macro F1 score over a classifier that used embeddings from a model pre-trained only on ImageNet. The smallest training dataset size of exhibited the largest improvement in macro F1 score, with the BYOL pre-trained model achieving a macro F1 score of % compared to % for the ImageNet pre-trained model, an improvement of %. The embedding model where BYOL pre-training was initialized with ImageNet weights performed marginally worse than the model that began BYOL pre-training with random weights: the average increase in macro F1 score over the ImageNet pre-trained model was % for the BYOL model with random initialization, compared to % for the BYOL model with ImageNet initialization.
While the linear probing results show vast improvements in representation quality when using BYOL pre-training, fine-tuning these pre-trained encoders using various semantic segmentation architectures resulted in different performance trends. Figure 8 shows that while the BYOL models that were initialized with ImageNet weights prior to BYOL pre-training outperformed ImageNet models with no additional pre-training, the BYOL models that were initialized with random weights yielded more variable performance depending on the specifics of the semantic segmentation architecture. The BYOL encoder with random initialization outperformed the ImageNet pre-trained encoder in terms of mean macro F1 score across the UPerNet and DeepLabV3+ models at , but lagged behind all other architectures at and . Overall, the ImageNet pre-trained models had a mean macro F1 score of %, while the BYOL model with random initialization had a mean macro F1 score of %. However, the BYOL model initialized with ImageNet weights prior to pre-training consistently outperformed the ImageNet pre-trained model across all architectures and training dataset sizes, achieving a mean macro F1 score of %.
Increasing the training dataset size generally led to improvements in macro F1 score across all architectures and pre-training strategies, though the magnitude of these improvements diminished as the dataset size increased. For example, the average increase in macro F1 score when increasing the training dataset size from to was %, while the average increase from to was only %. The impact of increasing training dataset size was most pronounced for the ImageNet pre-trained models, which saw a mean increase of % in macro F1 score when increasing the training dataset size from to , compared to only % for the BYOL (ImageNet initialization) models and % for the BYOL (random initialization) models. In terms of semantic segmentation architecture, the U-Net models had the highest mean macro F1 score across all training dataset sizes and pre-training strategies (71.20%), followed by Attention U-Net (70.85%), UPerNet (69.01%), DeepLabV3+ (67.96%), PAN (67.32%), and FCN (65.22%). No correlation between model size (in terms of number of parameters) and macro F1 score was observed (), indicating that there was no clear relationship between model complexity and performance in this context.
The best-performing pre-training, semantic segmentation architecture, and training dataset size combination based on mean macro F1 score across all k-folds was the BYOL model initialized with ImageNet weights used as the backbone in a U-Net architecture trained with samples, which achieved a mean macro F1 score of % and a mean overall accuracy of %, a % and % improvement over the U-Net model using the same training dataset size but with an ImageNet pre-trained backbone, respectively. To illustrate the qualitative benefits of the BYOL pre-training approach, we highlight a particularly strong individual model run. Run #2 of the U-Net model with the BYOL backbone (ImageNet initialization) fine-tuned using samples achieved a macro F1 score of %. A qualitative assessment of this model’s inference results is shown in Figure 9, demonstrating that not only did the model outperform its ImageNet-only counterpart quantitatively (an increase of % macro F1 score for the same data split), but it also produced land cover maps with sharper object boundaries and fewer classification artifacts. Samples (a), (b), and (c) in Figure 9 show overall improvements in classification accuracy with additional pre-training across multiple land cover classes, particularly in homogeneous regions such as water bodies and fields (though some misclassifications still occur, such as in (c)). The importance of object localization is highlighted in samples (d), (e), and (f), where the ImageNet-only model tends to yield acceptable overall classification accuracy, but fails to accurately delineate small features such as roads and buildings, which are better captured by the model with additional BYOL pre-training. While these examples are illustrative and do not change the overall ranking reported above, they highlight the qualitative improvements of the BYOL pre-training approach, which are particularly important given the spatial resolution of the ground truth labels.
Despite this specific model’s large individual performance, it was an outlier in terms of performance across all k-folds, as the mean macro F1 score and overall accuracy for this model were % and %, respectively, falling short of the best-performing configuration in terms of mean macro F1 score. As such, the U-Net model with a BYOL backbone initialized with ImageNet weights and fine-tuned using samples was considered the best-performing model overall and was used to generate the final land cover map for the entire state of Mississippi. The 4 models trained using this best-performing configuration as part of the k-fold cross-validation procedure were subsequently used to generate a final land cover map for the entire state of Mississippi using the model ensembling and test-time augmentation procedure outlined in Section 3.4. A visualization of the final comprehensive resolution land cover map for Mississippi is shown in Figure 10. We note the presence of a large vertical stripe of barren land in the Mississippi Delta region that stands in contrast to the surrounding cropland; this is an artifact of a data acquisition anomaly in the 2023 NAIP imagery, where a portion of the imagery was collected during a period when fields were fallow and thus appeared barren. This highlights a unique challenge of mapping agricultural activity using imagery acquired across multiple dates, as phenological changes in cropland can cause inconsistent land cover classifications. Still, despite this anomaly, our final land cover product offers a complete and precise spatial inventory of natural resources and human developments across the entire state. For example, water bodies appear to be delineated with high precision, revealing the presence of small streams and ponds, forested areas are well captured with fragmentation appropriate to the landscape context (e.g., trees interspersed within urban developments), and developed areas with small buildings and road networks are distinct and mapped with adequate spatial detail.
The confusion matrix from final accuracy assessment is shown in Table 3, demonstrating that our model yields high accuracy when classifying natural resources, while struggling with land cover types that are more anthropogenically influenced. For example, the barren land class has a moderately high user’s accuracy of %, but a low producer’s accuracy of %, and the model has a tendency to misclassify barren land pixels as impervious surfaces, herbaceous, and cultivated crops. The cultivated crops class suffers from the opposite problem, with a high producer’s accuracy of % but a low user’s accuracy of %, with barren land and herbaceous classes being frequently misclassified as cultivated crops. On the other hand, the open water and tree canopy/woody vegetation classes are classified with high accuracy, with F1 scores of % and %, respectively, highlighting the model’s strength in identifying these natural features. Overall, the final product’s macro F1 score and overall accuracy were % and %, respectively, while the mean class-wise producer’s accuracy and user’s accuracy were % and %, respectively. We do note that the large amount of forested land in Mississippi makes a balanced assessment challenging, as forested land accounted for over % of all sampled assessment points, making the overall accuracy metric less informative due to the class imbalance: we thus emphasize the macro F1 score and individual class-wise metrics as a more holistic measure of model performance in this context.
| Open | |||||||||
| Water | Impervious | ||||||||
| Structures | Impervious | ||||||||
| Surfaces | Barren | ||||||||
| Land | Tree Canopy/ | ||||||||
| Woody Vegetation | Low Vegetation/ | ||||||||
| Herbaceous | Cultivated | ||||||||
| Crops | Unclassified | Producer’s | |||||||
| Accuracy | |||||||||
| Open Water | 511 | 0 | 10 | 11 | 18 | 0 | 21 | 89.34% | |
| Impervious | |||||||||
| Structures | 0 | 15 | 3 | 3 | 5 | 0 | 0 | 75.47% | |
| Impervious | |||||||||
| Surfaces | 1 | 183 | 17 | 1 | 22 | 0 | 1 | 79.57% | |
| Barren Land | 6 | 68 | 944 | 22 | 451 | 144 | 5 | 57.42% | |
| Tree Canopy/ | |||||||||
| Woody Vegetation | 1 | 2 | 6 | 14712 | 366 | 6 | 22 | 97.32% | |
| Low Vegetation/ | |||||||||
| Herbaceous | 7 | 20 | 138 | 376 | 3699 | 492 | 17 | 77.86% | |
| Cultivated Crops | 0 | 0 | 17 | 2 | 38 | 1243 | 1 | 95.54% | |
| Unclassified | 2 | 6 | 7 | 753 | 86 | 1 | 413 | 32.29% | |
| User’s Accuracy | 96.78% | 62.24% | 82.66% | 92.64% | 78.95% | 65.91% | 86.04% | Overall Accuracy | |
| 87.14% | |||||||||
Figure 11 provides a visual comparison of the final land cover product against comparable existing land cover datasets (NLCD, ESRI Annual Land Use Land Cover [104]) over various land cover settings in Mississippi, illustrating the increase in spatial detail our resolution product provides over existing products with coarser spatial resolutions ( for ESRI LULC and for NLCD). We observe that our product is more suitable for applications requiring fine spatial detail, high spatial accuracy, and precise object boundaries given the improved spatial resolution over these existing products. For example, agricultural fields are precisely delineated in our product, while they appear as coarse blocks of land cover in the NLCD and ESRI LULC products. In urban areas, the mix of natural and developed land cover types interspersed within residential neighborhoods is well captured in our product, while the coarser products tend to generalize these areas into a single land cover type. In forested areas, small clearings, roads, and barren patches are captured in our product, while they are often absent in the coarser products. However, these existing products are derived from satellite imagery composites and represent the land cover over the span of a given year, while our product is derived from aerial imagery collected during a specific acquisition window in 2023. As observed in the Mississippi Delta region, this can lead to discrepancies in land cover classification due to phenological changes in agricultural land cover over time, which is a limitation of our product compared to these existing products that integrate data over longer time periods.
The final land cover product’s macro F1 score represents an increase of % over the mean performance of the base k-fold models, which can be attributed to the model ensembling and test-time augmentation procedures outlined in Section 3.4. Table 4 provides a class-wise comparison of the F1 scores between the performance of the models in the k-fold cross-validation and their performance when ensembled. Notably, the impervious surfaces and barren land classes saw dramatic improvements in F1 score after ensembling, with increases of % and %, respectively. Other classes saw more modest improvements, while the unclassified class saw a decrease of % in F1 score after ensembling. The use of the model ensemble approximately quadrupled the time needed for inference compared to a single model, highlighting the trade-off between accuracy and computational efficiency when using model ensembling and test-time augmentation. For our purposes, we deemed the increase in accuracy worth the additional computational cost, which was not prohibitive given the available computational resources, and elected to use the ensembled predictions for the final land cover product.
| Open | ||||||||
| Water | Impervious | |||||||
| Structures | Impervious | |||||||
| Surfaces | Barren | |||||||
| Land | Tree Canopy/ | |||||||
| Woody Vegetation | Low Vegetation/ | |||||||
| Herbaceous | Cultivated | |||||||
| Crops | Unclassified | |||||||
| Mean Ensemble F1 | 92.24 | 63.39 | 63.96 | 94.84 | 77.01 | 76.55 | 49.04 | |
| Final F1 | 92.91 | 69.85 | 67.77 | 94.93 | 78.40 | 78.00 | 46.96 | |
| Improvement | +0.67 | +6.46 | +3.81 | +0.09 | +1.40 | +1.45 | -2.09 |
5 Discussion
This study demonstrated that deep learning semantic segmentation models for dense land cover classification at VHSR can be effectively trained with limited labeled data, and that the resulting models can be used to produce highly detailed land cover maps over large areas. As shown in Figure 11, the final land cover product produced using our approach provides a detailed characterization of surface phenomena across the entire state of Mississippi with improved spatial fidelity compared to existing data products such as NLCD data and data from the ESRI annual land cover dataset [104]. Our land cover product captures fine-scale land cover features such as individual buildings, roads, small bodies of water, precise agricultural field delineations, forest clearings, and mixed urban-vegetation interfaces that are indistinguishable in the NLCD land cover data. As Boyle et al. [105] and Fisher et al. [14] noted, the ability to classify fine-scale land cover features is critical for a wide variety of applications, and has been shown to have a substantial impact on the accuracy and utility of downstream applications that rely on land cover data, such as water quality modeling, land use assessment, and conservation biology. However, the 8 land cover classes used in our product are more general than those used in the NLCD dataset, which can limit the expressiveness of our product in certain contexts (e.g., distinguishing between different forest types, wetlands, developed intensities, etc.). Additionally, confusion between certain classes is evident in the thematic accuracy assessment results shown in Table 3, demonstrating that while overall accuracy is high, certain classes remain challenging to classify accurately. For example, there is substantial confusion between the barren land and impervious surfaces classes, which is likely due to the fact that these surfaces are spectrally and texturally similar in the NAIP imagery used in this work. Still, given the high spatial resolution and novelty of our data product, we believe it will be highly suitable for downstream work that requires highly detailed information of surface features at both small and large scales.
Core to the success of our approach is the use of self-supervised learning to pre-train an image encoder using large amounts of unannotated NAIP imagery prior to fine-tuning on the target land cover classification task. We achieved high land cover classification accuracy with only 1,000 labeled training samples, which is a small fraction of the amount of labeled data typically used to train deep learning models for similar tasks. For example, the ISPRS Potsdam and Vaihingen datasets can be used to create 23,750 and 8,873 256256 pixel training patches, respectively [40, 41]. Additionally, these datasets only cover small urban areas, while our work concerns a substantially larger extent with multiple diverse ecoregions. The results of our experiment provide strong evidence to support this claim, and we applied this methodology to produce a resolution land cover map for the entire state of Mississippi. We note that for linear probing, the embeddings produced by the encoder pre-trained using BYOL alone performed the best, slightly outperforming the BYOL model initialized with ImageNet weights and substantially outperforming the ImageNet pre-trained model. One observation from the quantitative assessment of the model is the discrepancy between the performance of the pre-trained models when used as fixed feature extractors for linear probing compared to when they are fine-tuned as part of an end-to-end trainable semantic segmentation architecture. When only working with fixed embeddings, the BYOL model with random initialization outperformed the BYOL model initialized with ImageNet weights, which in turn outperformed the ImageNet pre-trained model. This suggests that the embeddings produced by the BYOL model with random initialization are somewhat linearly separable with respect to the land cover classes of interest, even in the absence of any supervised fine-tuning. This is a promising result, as it contributes to the growing body of evidence that self-supervised learning methods can be just as effective (if not more so) than traditional supervised pre-training methods for remote sensing applications [106, 107, 108]. These results also provided a quantitative demonstration of the domain gap between the natural, general-purpose images found in ImageNet and remote sensing imagery, and the benefits of pre-training on in-domain data, as echoed by findings in other recent works [109]. However, when fine-tuning the entire model as part of a semantic segmentation architecture, the BYOL model initialized with ImageNet weights consistently outperformed both the BYOL model with random initialization and the ImageNet pre-trained model across all architectures and training dataset sizes. This is a critical observation, as it indicates that the broader, more general image features learned from ImageNet are potentially relevant to the land cover classification task when a more complex decoder network is used to learn task-specific features during fine-tuning, suggesting that in spite of the apparent domain gap, ImageNet pre-training is still worthwhile for remote sensing applications.
A key feature of the BYOL approach used in this study is that it is model-agnostic in nature, meaning that it can theoretically be applied to any CNN or Transformer-based image encoder architecture, and these encoders can then be transferred into any downstream architecture for fine-tuning. On the other hand, many approaches to foundation model pre-training for remote sensing applications currently rely on ViT-based architectures [110, 111, 112], which have not yet seen wide adoption for dense prediction tasks in remote sensing due to the computational complexity associated with training and inference using such models, while CNN-based semantic segmentation architectures are still widely used and well understood in the remote sensing community [113, 114]. Additionally, given recent results in computer vision research that show that CNNs can achieve comparable performance to Transformers on a variety of computer vision tasks when designed appropriately [115, 116], it is still unclear whether Transformer-based architectures are strictly superior to CNNs for remote sensing applications across the board. Thus, the model-agnostic nature of BYOL is a significant advantage, as it allows for flexibility in choosing the most appropriate architecture for the target task without being constrained to a specific model type. For example, we transferred the BYOL pre-trained ResNet-101 encoder into multiple semantic segmentation architectures (FCN, U-Net, Attention U-Net, DeepLabV3+, UPerNet, PAN) and found that using the BYOL pre-trained encoder consistently improved performance across all architectures compared to using an ImageNet pre-trained encoder alone. Given the wide availability of unannotated VHSR imagery from sources such as NAIP and government satellite programs, we believe that self-supervised learning methods such as BYOL provide a promising avenue for future research in remote sensing applications, where labeled data is often scarce but unannotated data is abundant.
We note that while our results demonstrate the effectiveness of BYOL for pre-training an image encoder for land cover classification at VHSR, other self-supervised learning methods may also be effective for this task. One notable restriction of BYOL and similar transformation-based self-supervised learning methods is that they tend to rely on image augmentation routines that are specifically designed for three-channel RGB imagery, which can limit their applicability to multispectral or hyperspectral remote sensing imagery that contains more than three bands. For example, a central component of BYOL’s augmentation pipeline involves adjusting the brightness, contrast, saturation, and hue of an input image. The saturation and hue adjustment transformations are explicitly defined for three-channel imagery as a mapping from the RGB color space to the HSV color space, and thus cannot be directly applied to imagery with more than three bands. For remote sensing applications, this is a significant limitation of transformation-based self-supervised learning methods that rely on image augmentations to learn transformation-invariant image representations, restricting their applicability to only 3-band imagery. While we achieved strong results using BYOL on 3-band color-infrared NAIP imagery, 4 or more bands may be necessary to effectively utilize spectral information for other remote sensing applications, such as vegetation mapping, crop classification, or soil property estimation. Future work should continue to explore the implications of self-supervised learning for remote sensing applications under real-world constraints beyond academic benchmarks and compute-intensive techniques. Further, for label-scarce scenarios, little information on optimal sampling strategies for patch-based dense classification tasks is currently available in the literature, as opposed to more well-studied pixel-based sampling strategies for traditional machine learning classifiers [117, 75, 6].
6 Conclusion
In this work, we have demonstrated that accurate land cover classification at VHSR can be achieved using as few as 1,000 256256-pixel patches of labeled data by utilizing a BYOL self-supervised pre-training objective to learn meaningful image representations from large amounts of unlabeled VHSR imagery prior to fully supervised fine-tuning. We found that initializing a model with ImageNet weights and then continuing pre-training with BYOL was consistently the best-performing approach, suggesting that there is still value in utilizing models pre-trained on general-purpose imagery for remote sensing applications, even if there is a demonstrable domain gap. Still, pre-training a model using BYOL from a random initialization was often competitive with ImageNet pre-training alone, demonstrating the power of in-domain self-supervision and the need for remote sensing-specific pre-training, even beyond the context of foundation models. We used this approach to produce a resolution land cover map for the entire state of Mississippi, USA, which captures fine-scale land cover features that are indistinguishable in existing land cover products such as the NLCD. This achievement serves as a blueprint for both researchers applying geospatial deep learning models to data-scarce scenarios as well as practitioners seeking to produce high-quality land cover maps at VHSR over major geographic extents with limited human resources for collecting ground truth data.
Acknowledgments
This work was supported by the Mississippi Space Grant Consortium (MSSGC) Graduate Fellowship Program through NASA funding, as well as by the Mississippi Agricultural and Forestry Experiment Station (MAFES) Special Research Initiative. Dr. Martins and Lima were supported by the NASA Early Career Investigator Program in Earth Science (award #80NSSC24K1040). We also wish to acknowledge the University of Mississippi’s Mississippi Center for Supercomputing Research and Mississippi State University’s High Performance Computing Collaboratory for providing the computational resources necessary to complete this work. An online dashboard to visualize and interact with the final land cover product can be found at https://www.gcerlab.com/projects/msland, while the final land cover product can be downloaded from https://doi.org/10.5281/zenodo.16762344. The ground truth patches used to train the model can be found at https://doi.org/10.5281/zenodo.15670823.
CRediT authorship contribution statement
Dakota Hester: Methodology, Software, Formal analysis, Data curation, Visualization, Writing - original draft. Vitor S. Martins: Conceptualization, Methodology, Supervision, Writing - review & editing, Funding acquisition, Project administration. Lucas B. Ferreira: Methodology, Writing - review & editing. Thainara M. A. Lima: Writing - review & editing.